tickets to a local movie were sold at $6.00 for adults & $4.50 for students. If 270 tickets were sold for a total of $1590.00, how many student tickets were sold?

Answer by ewatrrr(24785)


Hi,
270 tickets were sold
Let x and (270-x) represent the number of adult and student tickets respectively
Question States***
$6*x + $4.50(270-x) = $1590
solving for x
6x + 4.50*270 - 4.50x = 1590
1.50x = 1590 - 1215
1.50x = 375
x = 250, the number of adult tickets. Student tickets = 20
CHECKING our Answer***
$6*250 + 20*$4.50 = $1590